# Dynamics of quantum Fisher and Wigner-Yanase skew information following a noisy quench

### v1

@@ VI Conclusions
Motivated by the fundamental questions of how quantum correlations are generated under noisy quenches across quantum critical points and by the ubiquity of control‐field fluctuations in realistic implementations, we have examined the dynamics of quantum Fisher information (QFI) and Wigner–Yanase skew information (WYSI) in Ising chain subjected to noiseless and noisy linear driven transverse field. In the absence of noise, and very large ramp time scale, the critical points $h_{c}=\pm 1$ are signaled by both QFI and ($x$) component of WYSI where become minimum. while, as the ramp time scale decreases, both QFI and all component of WYSI represent cusps and the critical points. Moreover, both QFI and WYSI grow monotonically with the ramp time scale $\tau$, reflecting enhanced quantum coherence and metrological usefulness as the quench slows down, and saturate to their static (adiabatic) values for large $\tau$. This result is contrary to the behavior of entanglement between next-nearest-neighbor spins, which reveals suppression in accordance with $\tau^{-1/2}$ [44, 45, 98]. When a Gaussian white noise of strength $\xi$ is added to the control field, the dynamics change qualitatively: both QFI and WYSI no longer simply saturate to the static values by increasing the ramp time scale, but instead develop a maximum at an intermediate ramp time scale $\tau_{m}$, beyond which their are suppressed by dephasing. By analyzing the logarithm of QFI and WYSI versus $\tau$, we demonstrated an exponential decay of both measures with ramp time scale under noise. While the density of defects in the presence of the noise does not scale exponentially with the ramp time scale [96, 99]. Moreover, the $\tau_{m}$ that maximizes each QFI and WYSI curve scales as $\tau_{m}(\xi)\propto(\xi^{2})^{\delta},$ with a universal exponent $\delta\approx 2/3\,$ for both nearest‐ and next‐nearest‐neighbour spins. This exponent corresponds to the exponent that dictates the scaling of the optimal time for achieving minimal defect production in a standard quantum annealing scheme. Whether the identity of the exponents, both detected in quenches that emulate standard quantum annealing schemes, is merely coincidental or indicates a deeper link between the two phenomena—minimization of KZM defects and maximization of QFI and WYSI in the presence of noise—remains an open question.
@@ Acknowledgment
Appendix A Reduced density matrix and Concurrence
To calculate quantum correlations, we use the two-spin reduced density matrix $\varrho_{l,m}(t)$ for spins at the sites $\ell$ and $m=\ell+r$, which is given by [100, 25].
$\varrho_{\ell,m}(t)=\left(\begin{array}[]{cccc}\rho_{11}&0&0&\rho_{14}\\ 0&\rho_{22}&\rho_{23}&0\\ 0&\rho_{23}^{\ast}&\rho_{33}&0\\ \rho_{14}^{\ast}&0&0&\rho_{44}\\ \end{array}\right),$ (29)
where matrix elements can be written in terms of one- and two-point correlation functions [100, 25]
$\rho_{11}$ $=\left\langle M^{z}_{\ell}\right\rangle+\left\langle s^{z}_{\ell}s^{z}_{m}\right\rangle+\frac{1}{4};\quad\rho_{22}=\rho_{33}=-\left\langle s^{z}_{\ell}s^{z}_{m}\right\rangle+\frac{1}{4},$ (30)
$\rho_{44}$ $=-\left\langle M^{z}_{\ell}\right\rangle+\left\langle s^{z}_{\ell}s^{z}_{m}\right\rangle+\frac{1}{4},$
@@ Appendix A Reduced density matrix and Concurrence
It should be mentioned that, while in the time independent magnetic field $h_{0}(t)=h$ the two-point spin correlations $\langle s_{\ell}^{x}s_{m}^{y}\rangle$ and $\langle s_{\ell}^{y}s_{m}^{x}\rangle$ are zero, they are generated under quench of the magnetic field. Using the Jordan-Wigner transformation it has been shown that two points spin-spin correlation functions can be expressed as [100, 25]
$S_{r}^{xx}=\langle s_{1}^{x}s_{1+r}^{x}\rangle$ $=\frac{1}{4}\langle B_{1}A_{2}B_{2}\ldots A_{r}B_{r}A_{r+1}\rangle,$ (31)
$S_{r}^{yy}=\langle s_{1}^{y}s_{1+r}^{y}\rangle$ $=\frac{(-1)^{n}}{4}\langle A_{1}B_{2}A_{2}\ldots B_{r}A_{r}B_{r+1}\rangle,$
$S_{r}^{zz}=\langle s_{1}^{z}s_{1+r}^{z}\rangle$ $=\frac{1}{4}\langle A_{1}B_{1}A_{r+1}B_{r+1}\rangle,$
$S_{r}^{xy}=\langle s_{1}^{x}s_{1+r}^{y}\rangle$ $=-i\frac{1}{4}\langle B_{1}A_{2}B_{2}\ldots A_{r}B_{r}B_{r+1}\rangle,$
$S_{r}^{yx}=\langle s_{1}^{y}s_{1+r}^{x}\rangle$ $=i\frac{(-1)^{r}}{4}\langle A_{1}B_{2}A_{2}\ldots B_{r}A_{r}A_{r+1}\rangle.$
@@
$A_{\ell}=c_{\ell}^{\dagger}+c_{\ell};\quad B_{\ell}=c_{\ell}^{\dagger}-c_{\ell}$
[…]
@@
Since the model is exactly solvable and can be written as the sum of $N/2$ non-interacting terms in momentum space. Therefore the four types of two-fermion correlation functions in the real space (Eq. (32)) can be calculated easily by transforming to the momentum space, i.e.,
$\langle c_{\ell}^{\dagger}c_{m}^{\dagger}\rangle=$ $\frac{1}{N}\sum_{p,q}e^{-i(p{\ell}+qm)}\langle\psi_{k}(t)|c_{p}^{\dagger}c_{q}^{\dagger}|\psi_{k}(t)\rangle$ (33)
$=$ $\frac{1}{N}\sum_{k}e^{-ik(m-{\ell})}\Big{[}i\Big{(}|u_{k}(t)|^{2}-|v_{k}|^{2}(t)\Big{)}\frac{\sin(\theta_{k})}{2}$
$-\Big{(}u_{k}^{\ast}(t)v_{k}(t)\cos^{2}(\frac{\theta_{k}}{2})+u_{k}(t)v_{k}^{\ast}(t)\sin^{2}(\frac{\theta_{k}}{2})\Big{)}\Big{]},$
and
$\langle c_{\ell}c_{m}\rangle=$ $\frac{1}{N}\sum_{p,q}e^{i(p{\ell}+qm)}\langle\psi_{k}(t)|c_{p}c_{q}|\psi_{k}(t)\rangle$ (34)
$=$ $\frac{1}{N}\sum_{k}e^{ik(m-{\ell})}\Big{[}i\Big{(}|u_{k}(t)|^{2}-|v_{k}(t)|^{2}\Big{)}\frac{\sin(\theta_{k})}{2}$
$+\Big{(}u_{k}^{\ast}(t)v_{k}(t)\sin^{2}(\frac{\theta_{k}}{2})+u_{k}(t)v_{k}^{\ast}(t)\cos^{2}(\frac{\theta_{k}}{2})\Big{)}\Big{]},$
and analogous expressions for $\langle c_{\ell}^{\dagger}c_{m}\rangle$ is given by
$\langle c_{\ell}^{\dagger}c_{m}\rangle=$ $\frac{1}{N}\sum_{p,q}e^{-i(p{\ell}-qm)}\langle\psi_{k}(t)|c_{p}^{\dagger}c_{q}|\psi_{k}(t)\rangle$ (35)
$=$ $\frac{1}{N}\sum_{k}\Big{[}\!\Big{(}|v_{k}(t)|^{2}\sin^{2}(\frac{\theta_{k}}{2})+|u_{k}(t)|^{2}\cos^{2}(\frac{\theta_{k}}{2})\Big{)}$
$-i\Big{(}u_{k}^{\ast}(t)v_{k}-u_{k}(t)v_{k}^{\ast}(t)\Big{)}\frac{\sin(\theta_{k})}{2}\Big{]}e^{ik(m-{\ell})}.$
Using $|u_{-k}|^{2}=|u_{k}|^{2}$, $|v_{-k}|^{2}=|v_{k}|^{2}$, $v_{-k}u^{\ast}_{-k}=-v_{k}u^{\ast}_{k}$, $\theta_{-k}=-\theta_{k}$ and ${\ell}-m=r$, the two-points fermion operators correlations reduces to
$\langle{{A_{\ell}}{A_{{\ell}+r}}}\rangle=\frac{2i}{N}\sum_{k>0}\Big{[}\sin(kr){\rm Re}[(v_{k}(t)u_{k}^{\ast}(t)]\Big{]}+\delta_{r,0},$ (36)
$\langle{{B_{\ell}}{B_{{\ell}+r}}}\rangle=\frac{2i}{N}\sum_{k>0}\Big{[}\sin(kr){\rm Re}[v_{k}(t)u_{k}^{\ast}(t)]\Big{]}-\delta_{r,0},$
$\langle A_{{\ell}}B_{{\ell}+r}\rangle=$
$\frac{2}{N}\sum_{k>0}\Big{\{}(1-2|u_{k}(t)|^{2})\Big{[}\!\cos(kr)\cos(\theta_{k})-\sin(kr)\sin(\theta_{k})\Big{]}$
$-2{\rm Im}[v_{k}(t)u_{k}^{\ast}(t)]\Big{(}\sin(kr)\cos(\theta_{k})+\sin(kr)\cos(\theta_{k})\Big{]}\Big{\}}.$
where $\delta_{r,0}$ is kronecker delta. It should be mention that, the above two-points fermion operators correlation reduces to that of equilibrium case [101] if we take $u_{k}(t)=0$,
$\langle{{A_{\ell}}{A_{{\ell}+r}}}\rangle=$ $\delta_{r,0};\quad\langle{{B_{\ell}}{B_{{\ell}+r}}}\rangle=-\delta_{r,0},$ (37)
$\langle A_{{\ell}}B_{{\ell}+r}\rangle=$ $-\langle B_{{\ell}+r}A_{{\ell}}\rangle$
$=$ $\frac{1}{N}\sum_{k}\Big{[}\cos(kr)\cos(\theta_{k})-\sin(kr)\sin(\theta_{k})\Big{]},$
which, in the thermodynamic limit ($N\to\infty$), reduce to the expressions given in Ref. [101]
$\langle{{A_{\ell}}{A_{{\ell}+r}}}\rangle=$ $-\langle{{B_{\ell}}{B_{{\ell}+r}}}\rangle=\delta(r),$ (38)
$\langle A_{{\ell}}B_{{\ell}+r}\rangle=$ $-\langle B_{{\ell}+r}A_{{\ell}}\rangle$
[…]
In addition, for a quench, from $h({t}_{i})\!\rightarrow-\infty$ to $h(t_{f})\!\rightarrow+\infty$, the off-diagonal term of the density matrix $v_{k}(t)u_{k}^{\ast}(t)=(v^{\ast}_{k}(t)u_{k}(t))^{\ast}$ and $\theta_{k}$ become zero and the above two-points fermion operators correlation reduces to
$\langle{{A_{\ell}}{A_{{\ell}+r}}}\rangle=$ $-\langle{{B_{\ell}}{B_{{\ell}+r}}}\rangle=\delta_{r,0};$ (39)
$\langle A_{{\ell}}B_{{\ell}+r}\rangle=$ $-\langle B_{{\ell}+r}A_{{\ell}}\rangle=\frac{1}{N}\sum_{k}[1-2p_{k}(\tau_{k,f})]\cos(kr),$
which, in the thermodynamics limit ($N\rightarrow\infty$), are equivalent to
$\langle{{A_{\ell}}{A_{{\ell}+r}}}\rangle=$ $-\langle{{B_{\ell}}{B_{{\ell}+r}}}\rangle=\delta(r),$ (40)
$\langle A_{{\ell}}B_{{\ell}+r}\rangle=$ $-\langle B_{{\ell}+r}A_{{\ell}}\rangle$
@@
where $p_{k}(\tau_{k,f})=e^{-\pi\gamma^{2}\sin^{2}(k)\tau}$ [45, 40, 44].

### v3

@@ VI Conclusions
Motivated by the fundamental questions of how quantum correlations are generated under noisy quenches across quantum critical points and by the ubiquity of control‐field fluctuations in realistic implementations, we have examined the dynamics of quantum Fisher information (QFI) and Wigner–Yanase skew information (WYSI) in Ising chain subjected to noiseless and noisy linear driven transverse field. In the absence of noise, and very large ramp time scale, the critical points $h_{c}=\pm 1$ are signaled by both QFI and ($x$) component of WYSI where become minimum. While, as the ramp time scale decreases, both QFI and all component of WYSI represent cusps and the critical points. Moreover, both QFI and WYSI grow monotonically with the ramp time scale $\tau$, reflecting enhanced quantum coherence and metrological usefulness as the quench slows down, and saturate to their static (adiabatic) values for large $\tau$. This result is contrary to the behavior of entanglement between next-nearest-neighbor spins, which reveals suppression in accordance with $\tau^{-1/2}$ [47, 48, 102]. When a Gaussian white noise of strength $\xi$ is added to the control field, the dynamics change qualitatively: both QFI and WYSI no longer simply saturate to the static values by increasing the ramp time scale, but instead develop a maximum at an intermediate ramp time scale $\tau_{m}$, beyond which they are suppressed by dephasing. By analyzing the logarithm of QFI and WYSI versus $\tau$, we demonstrated an exponential decay of both measures with ramp time scale under noise. While the density of defects in the presence of the noise does not scale exponentially with the ramp time scale [100, 103]. Moreover, the $\tau_{m}$ that maximizes each QFI and WYSI curve scales as $\tau_{m}(\xi)\propto(\xi^{2})^{\delta},$ with a universal exponent $\delta\approx 2/3\,$ for both nearest‐ and next‐nearest‐neighbour spins. This exponent corresponds to the exponent that dictates the scaling of the optimal time for achieving minimal defect production in a standard quantum annealing scheme. Whether the identity of the exponents, both detected in quenches that emulate standard quantum annealing schemes, is merely coincidental or indicates a deeper link between the two phenomena, minimization of KZM defects and maximization of QFI and WYSI in the presence of noise, remains an open question. Moreover, in the presence of noise, the lack of an appropriate analytical solution corresponding to numerically challenging problems limits the ability to link the scaling exponent to the well-known critical exponents associated with equilibrium phase transitions.
@@ Acknowledgment
Appendix A Reduced Density Matrix, Fermionic Correlation Functions, and Concurrence
To evaluate quantum correlations, we employ the two-spin reduced density matrix $\varrho_{\ell,m}(t)$ for spins located at sites $\ell$ and $m=\ell+r$, which takes the form [104, 28]
$\varrho_{\ell,m}(t)=\begin{pmatrix}\rho_{11}&0&0&\rho_{14}\\ 0&\rho_{22}&\rho_{23}&0\\ 0&\rho_{23}^{\ast}&\rho_{33}&0\\ \rho_{14}^{\ast}&0&0&\rho_{44}\end{pmatrix}$, (30)
where the matrix elements are expressed in terms of one- and two-point correlation functions [104, 28].
$\rho_{11}$ $=\left\langle M^{z}_{\ell}\right\rangle+\left\langle s^{z}_{\ell}s^{z}_{m}\right\rangle+\tfrac{1}{4};\quad\rho_{22}=\rho_{33}=-\left\langle s^{z}_{\ell}s^{z}_{m}\right\rangle+\tfrac{1}{4},$ (31)
$\rho_{44}$ $=-\left\langle M^{z}_{\ell}\right\rangle+\left\langle s^{z}_{\ell}s^{z}_{m}\right\rangle+\tfrac{1}{4},$
@@ Appendix A Reduced Density Matrix, Fermionic Correlation Functions, and Concurrence
Note that in a time-independent magnetic field $h_{0}(t)=h$, the two-point spin correlations $\langle s_{\ell}^{x}s_{m}^{y}\rangle$ and $\langle s_{\ell}^{y}s_{m}^{x}\rangle$ vanish. However, they are dynamically generated during a quench of the magnetic field. Using the Jordan–Wigner transformation, it has been shown that the two-point spin–spin correlation functions can be written as [104, 28]
$S_{r}^{xx}=\langle s_{1}^{x}s_{1+r}^{x}\rangle$ $=\tfrac{1}{4}\langle B_{1}A_{2}B_{2}\ldots A_{r}B_{r}A_{r+1}\rangle,$ (32)
$S_{r}^{yy}=\langle s_{1}^{y}s_{1+r}^{y}\rangle$ $=\tfrac{(-1)^{n}}{4}\langle A_{1}B_{2}A_{2}\ldots B_{r}A_{r}B_{r+1}\rangle,$
$S_{r}^{zz}=\langle s_{1}^{z}s_{1+r}^{z}\rangle$ $=\tfrac{1}{4}\langle A_{1}B_{1}A_{r+1}B_{r+1}\rangle,$
$S_{r}^{xy}=\langle s_{1}^{x}s_{1+r}^{y}\rangle$ $=-\tfrac{i}{4}\langle B_{1}A_{2}B_{2}\ldots A_{r}B_{r}B_{r+1}\rangle,$
$S_{r}^{yx}=\langle s_{1}^{y}s_{1+r}^{x}\rangle$ $=\tfrac{i(-1)^{r}}{4}\langle A_{1}B_{2}A_{2}\ldots B_{r}A_{r}A_{r+1}\rangle.$
@@
$A_{\ell}=c_{\ell}^{\dagger}+c_{\ell},\quad B_{\ell}=c_{\ell}^{\dagger}-c_{\ell}.$ (33)
It is straightforward to show that
@@
Since the model is exactly solvable and can be written as a sum of $N/2$ non-interacting terms in momentum space, the four types of two-fermion correlation functions in real space [Eq. (34)] can be readily obtained by Fourier transforming to momentum space. For example,
$\langle c_{\ell}^{\dagger}c_{\ell+r}^{\dagger}\rangle=\mathrm{Tr}(\rho\,c_{\ell}^{\dagger}c_{\ell+r}^{\dagger})=\frac{1}{N}\sum_{p,q}e^{-i(p\ell+q(\ell+r))}\langle c_{p}^{\dagger}c_{q}^{\dagger}\rangle.$ (35)
We use the diagonal basis $\{|\phi_{k}^{\pm}\rangle\}_{k}$ of the mode Hamiltonians at fixed time $t$ to evaluate the trace in Eq. (35):
$\mathrm{Tr}[\rho_{k}c_{k}^{\dagger}c_{-k}^{\dagger}]=\langle\phi^{-}_{k}|\rho_{k}c_{k}^{\dagger}c_{-k}^{\dagger}|\phi^{-}_{k}\rangle+\langle\phi^{+}_{k}|\rho_{k}c_{k}^{\dagger}c_{-k}^{\dagger}|\phi^{+}_{k}\rangle$$\quad$ (36)
@@
where the superscript $(d)$ indicates that the density matrix is written in the diagonal basis. The matrix elements of $c_{k}^{\dagger}c_{-k}^{\dagger}$ can be computed from Eq. (4), leading to
$\langle c_{\l }^{\dagger}c_{\l +r}^{\dagger}\rangle=\frac{1}{N}\sum_{k>0}$ $\sin(kr)\Big[\sin(2\theta_{k})(\rho^{(d)}_{k,22}-\rho^{(d)}_{k,11})$ (38)
$+2i(\cos^{2}(\theta_{k})\,\rho^{(d)}_{k,12}+\sin^{2}(\theta_{k})\,\rho^{(d)}_{k,21})\Big$].
Similarly, one finds
[…]
$\langle c_{\l }^{\dagger}c_{\l +r}\rangle=$ $\frac{1}{N}\sum_{k>0}\cos(kr)\Big[2\cos^{2}(\theta_{k})\rho^{(d)}_{k,22}+2\sin^{2}(\theta_{k})\rho^{(d)}_{k,11}$ (40)
$\quad\quad\quad-i\sin(2\theta_{k})(\rho^{(d)}_{k,12}-\rho^{(d)}_{k,21})\Big$],
$\langle c_{\l }c_{\l +r}^{\dagger}\rangle\!=\!\frac{1}{N}\sum_{k>0}$ $\cos(kr)\Big[2\cos^{2}(\theta_{k})\rho^{(d)}_{k,11}+2\sin^{2}(\theta_{k})\rho^{(d)}_{k,22}$ (41)
$-i\sin(2\theta_{k})(\rho^{(d)}_{k,21}-\rho^{(d)}_{k,12})\Big$].
Substituting the results of Eqs. (38)–(41) into Eq. (34), we obtain the $A$- and $B$-type two-point functions that determine the spin correlations in Eq. (7). After simplification, the result reads
$\begin{aligned} &\langle A_{l}A_{l+r}\rangle=\frac{2i}{N}\sum_{k>0}\mathrm{Re}[\rho^{(d)}_{k,12}]\sin(kr)+\delta_{r,0},\\ &\langle B_{l}B_{l+r}\rangle=\frac{2i}{N}\sum_{k>0}\mathrm{Re}[\rho^{(d)}_{k,12}]\sin(kr)-\delta_{r,0},\\ &\langle A_{l}B_{l+r}\rangle=-\langle B_{l+r}A_{l}\rangle=\\ &\frac{1}{N}\sum_{k}\Big[(1-2\rho^{(d)}_{k,22})\big[\cos(kr)\cos(2\theta_{k})\!-\!\sin(kr)\sin(2\theta_{k})\big]\\ &\quad\quad-2\mathrm{Im}[\rho^{(d)}_{k,12}]\big(\cos(kr)\sin(2\theta_{k})+\sin(kr)\cos(2\theta_{k})\big)\Big].\end{aligned}$
We note that the two-point fermionic functions above reduce to their equilibrium forms if we set $\rho^{(d)}_{k,22}=\rho^{(d)}_{k,12}=\rho^{(d)}_{k,21}=0$. In this case, we obtain
$\begin{aligned} \langle A_{l}A_{l+r}\rangle=&\delta_{r,0};\quad\quad\langle B_{l}B_{l+r}\rangle=-\delta_{r,0},\\ \langle A_{l}B_{l+r}\rangle=&-\langle B_{l+r}A_{l}\rangle\\ =&\frac{1}{N}\sum_{k}\Big[\cos(kr)\cos(2\theta_{k})-\sin(kr)\sin(2\theta_{k})\Big].\end{aligned}$
In the thermodynamic limit $N\to\infty$, these correlation functions take the form
$\langle{{A_{l}}{A_{l+r}}}\rangle=\delta(r);\quad\quad\langle{{B_{l}}{B_{l+r}}}\rangle-\delta(r),$ (44)
$\langle A_{l}B_{l+r}\rangle=-\langle B_{l+r}A_{l}\rangle=\frac{1}{4\pi}\int_{0}^{2\pi}e^{ikr}e^{2i\theta_{k}}dk,\quad\quad$
which is consistent with the results reported in Ref. [105].
Furthermore, for a quench from $h(t_{\text{initial}})\to$-$\infty$ to $h(t_{\text{final}})\to+\infty$, the off-diagonal elements of the density matrix, $\rho^{(d)}_{k,12}(t)=\big(\rho^{(d)}_{k,21}(t)\big)^{\ast}$, vanish, as does $\theta_{k}(t)$. Consequently, the two-point fermionic functions in Eq. (A) reduce to
$\langle{{A_{l}}{A_{l+r}}}\rangle$ $=\delta_{r,0};\quad\quad\langle{{B_{l}}{B_{l+r}}}\rangle=-\delta_{r,0},$ (45)
$\langle A_{l}B_{l+r}\rangle$ $=-\langle B_{l+r}A_{l}\rangle=\frac{1}{N}\sum_{k}(1-2\rho^{(d)}_{k,22})\cos(kr),$
which, in the long-time thermodynamic limit, reproduces the results of Refs. [48, 43, 47],
$\langle{{A_{l}}{A_{l+r}}}\rangle=$ $\delta(r);\quad\quad\langle{{B_{l}}{B_{l+r}}}\rangle=-\delta(r),$ (46)
$\langle A_{l}B_{l+r}\rangle=$ $-\langle B_{l+r}A_{l}\rangle=\frac{1}{2\pi}\int_{\pi}^{\pi}[1-2p_{k}(t)]\cos(kr)dk$
$=$ $-\frac{2}{\pi}\int_{0}^{\pi}p_{k}(t)\cos(kr)dk,$
where
$\lim_{t\to\infty}\rho^{(d)}_{k,22}(t)=e^{-4\pi\sin^{2}(k)\,\tau}\equiv p_{k}(t),$
with $\tau$ denoting the time scale of the quench [48, 43, 47].